### v2

@@ 4.5 Training
Network training is performed with the AdamW optimizer loshchilov2017decoupled , combined with a decaying learning rate schedule. For the model with the subtraction operator, the learning rate decreases from $10^{-3}$ to $10^{-5}$ over 120 epochs. For the model with the dot-product operator, it decreases from $10^{-4}$ to $10^{-6}$ over 240 epochs.
@@ 5.3 PID Performance
The PID performance is further evaluated using the particle separation power defined in eq. 6. Figure 8 presents the $K/\pi$ separation $S_{K,\,\pi}$ as a function of track momentum. The tracks originate from the collision point with a polar angle of $60^{\circ}$, corresponding to a track length of approximately 1.4 m. For the dNN/dxx results (shown in red circulars), the conclusions are consistent with those in table 3: the GraphPT method significantly enhances PID performance compared to the traditional truncated mean approach, with the dot-product operator yielding the best overall $K/\pi$ separation. Specifically, the improvements range from approximately 10% (5%) to 20% (15%) in the momentum interval from 5 to 20 GeV/$c$ for the dot-product (subtraction) operator.
@@ 6 Conclusion and Outlook
Looking ahead, several aspects require further study. First, the GraphPT model itself still needs optimization in terms of network architecture and hyperparameters. Second, in this work the model was trained on a fix-angle sample; future studies should extend the training to much larger and more diverse datasets. This will also necessitate improving the computational efficiency of the model to enable large-scale training.
The ultimate goal of this study is to access the feasibility of PID performance using the dNN/dxx method. To this end, further validation against detector R&D is essential. For the CEPC TPC, two TPC prototypes have already been designed. Beam tests are planned to evaluate their performance and provide experimental validation of the high-granularity readout approach. One such beam test has scheduled at Deutsches Elektronen-Synchrotron (DESY) diener2019desy , within the LCTPC Collaboration and the CERN R&D Collaboration on gaseous detector (DRD1) colaleo2024drd1 .

### v3

@@ 4.5 Training
Network training is performed with the AdamW optimizer Loshchilov and Hutter (2017), combined with a decaying learning rate schedule. For the model with the subtraction operator, the learning rate decreases from $10^{-3}$ to $10^{-5}$ over 120 epochs. For the model with the dot-product operator, it decreases from $10^{-4}$ to $10^{-6}$ over 240 epochs.
@@ 5.3 PID Performance
The PID performance is further evaluated using the particle separation power defined in eq. 6. Figure 8 presents the $K/\pi$ separation power $S_{K,\,\pi}$ as a function of track momentum. The tracks originate from the collision point with a polar angle of $60^{\circ}$, corresponding to a track length of approximately 1.4 m. For the dNN/dxx results (shown in red circles), the conclusions are consistent with those in table 3: the GraphPT method significantly enhances PID performance compared to the traditional truncated mean approach, with the dot-product operator yielding the best overall $K/\pi$ separation. Specifically, the improvements range from approximately 10% (5%) to 20% (15%) in the momentum interval from 5 to 20 GeV/$c$ for the dot-product (subtraction) operator.
@@
To further assess the robustness of the GraphPT model, we apply it to a finer-granularity readout with a pad size of $200\times 200$ $\mu\textrm{m}^{2}$. An even larger performance gain is observed with the GraphPT model. Additional details are provided in appendix A.
@@ 6 Conclusion and Outlook
Looking ahead, several aspects require further study. First, the GraphPT model itself still needs optimization in terms of network architecture and hyperparameters. Second, in this work the model was trained on a fix-angle sample; future studies should extend the training to much larger and more diverse datasets. This will also necessitate improving the computational efficiency of the model to enable large-scale training. Moreover, as with any supervised-learning approach, the performance of the model depends on the consistency between the training data and the data it is applied to. Periodic retraining and calibration with real data will be essential to maintain robustness against systematic effects such as spacial inhomogeneities in detector response, time variations in TPC operating conditions, and field distortions induced by ion backflow.
The ultimate goal of this study is to access the feasibility of PID performance using the dNN/dxx method. To this end, further validation against real detectors is essential. For the CEPC TPC, two TPC prototypes have already been designed. Beam tests are planned to evaluate their performance and provide experimental validation of the high-granularity readout approach. One such beam test has scheduled at Deutsches Elektronen-Synchrotron (DESY) Diener et al. (2019), within the LCTPC Collaboration and the CERN R&D Collaboration on gaseous detector (DRD1) Colaleo et al. (2024).
@@ Acknowledgements.
Appendix A Validation Using a Reduced Pad Size of $200\times 200$ $\mu\textrm{m}^{2}$
To further validate the GraphPT model, an alternative simulation configuration is employed. In this setup, the detector is simulated with a smaller pad size of $200\times 200$ $\mu\textrm{m}^{2}$. As shown in figure 9, the increased granularity retains more of the underlying raw information: the hits become sparser, and both primary and secondary electrons are more readily detected. This richer but less compressed input makes the reconstruction task more challenging. In this simulation, only the pad size is modified; all other aspects of the detector and electronics design remain unchanged. Therefore, aside from the reduced pad size, all parameters are identical to those used in the default simulation.
The neural network architecture follows the design described in section 4.3. For the self‑attention block, only the dot‑product operator—–shown to deliver state‑of‑the‑art performance—–is used. The model is trained with a learning rate decreasing from $10^{-4}$ to $10^{-5}$ over 140 epochs. Following the procedure described in section 5.1, the threshold $\tau$ is optimized to $10^{-4}$; the fraction $\alpha$ and the number of combined pad rows used in the truncated mean method are optimized to 0.60 and 4, respectively.
Figure 9: Predicted pad classes for a track segment with a pad size of $200\times 200$ $\mu\textrm{m}^{2}$. (a) Pads predicted as negatives are marked by circles for the GraphPT and stars for the truncated mean method, respectively; all other pads are predicted as positives. The red box highlights the zoomed-in region. (b) The zoomed-in region using only the GraphPT method. (c) The zoomed-in region using only the truncated mean method.
Table 4: $K/\pi$ separation power obtained using the dNN/dxx method with the truncated mean (TM) and the GraphPT (ML) model, along with their relative difference, defined as |ML - TM|/TM.
Pad Length = 500 $\mu\textrm{m}$ Momentum (GeV/c)
5 10 15 20
Truncated Mean $5.21\pm 0.04$ $4.74\pm 0.04$ $4.00\pm 0.04$ $3.43\pm 0.03$
GraphPT (Dot-prod.) $6.43\pm 0.05$ $5.40\pm 0.04$ $4.44\pm 0.04$ $3.69\pm 0.03$
Relative Difference (%) $23.3\pm 1.4$ $4.0\pm 1.3 11$$1.0\pm 1.4$ $7.6\pm 1.4$
Pad Length = 200 $\mu\textrm{m}$ Momentum (GeV/c)
5 10 15 20
Truncated Mean $5.04\pm 0.04$ $4.60\pm 0.04$ $3.93\pm 0.04$ $3.42\pm 0.04$
GraphPT (Dot-prod.) $6.83\pm 0.05$ $5.90\pm 0.05$ $4.79\pm 0.04$ $3.94\pm 0.03$
Relative Difference (%) $35.5\pm 1.5$ $28.3\pm 1.5$ $21.7\pm 1.5$ $15.2\pm 1.5$
The classification results for a track segment are shown in figure 9. A conclusion similar to the $500\times 500$ $\mu\textrm{m}^{2}$ case can be drawn: the GraphPT model produces almost no false-negatives, whereas the truncated mean method is more aggressive and yields several false-negative classifications, thereby reducing signal efficiency. Table 4 summarizes the resulting $K/\pi$ separation power. Compared with the $500\times 500$ $\mu\textrm{m}^{2}$ configuration, the truncated mean performance is slightly degraded. This is mainly due to the reduced signal-to-noise ratio: although the noise level per pad remains $100~e^{-}$, the total number of pads increases by a factor of 6.25. Because the truncated mean method suppress secondary-electron contributions by removing entire row pads, it cannot fully exploit the high-granularity information. In contrast, the GraphPT model can perform pad-by-pad suppression, leading to a significantly improved $K/\pi$ separation power for the $200\times 200$ $\mu\textrm{m}^{2}$ readout. The improvement is approximately 35% to 15% across the momentum range from 5 to 20 GeV/$c$.
Code Availability Statement
This article has associted code in a code repository. The code for the GraphPT model is publicly available at https://github.com/littlepi/TpcDndxGraphPT.